OpenStudy (anonymous):

Help with this function. PLEASE. The function H is defined below h(x)=x+3/x^2-64 Find all the values of x that are not in the domain of h.

OpenStudy (ash2326):

\[h(x)=\frac{x+3}{x^2-64}\] @maddiemaze The values of x for which h(x) becomes undefined are not in the domain. Do you have any idea, how we should proceed?

OpenStudy (anonymous):

No

OpenStudy (ash2326):

OK, if you put x=8 , what would you get for h(x)?

OpenStudy (anonymous):

8h

OpenStudy (ash2326):

\[h(x)=\frac{x+3}{x^2-64}\] x=8 \[h(8)=\frac{8+3}{8^2-64}\] Can you try to solve the left side?

OpenStudy (anonymous):

11/0

OpenStudy (ash2326):

Good :) \[\frac{11}{0}=\infty\] Which is undefined, so if the denominator becomes 0, then h(x) would be undefined. @maddiemaze is there any other value of x for x^2-64=0??

OpenStudy (anonymous):

if its undefined than I dont think there would be another value

OpenStudy (ash2326):

Solve for \[x^2-64=0\] It's a quadratic, it'd have two values of x

OpenStudy (anonymous):

8

OpenStudy (ash2326):

and the other one?

OpenStudy (anonymous):

I have no idea

OpenStudy (ash2326):

\[x^2-64\] \[x^2-8^2=0\] We know \[a^2-b^2=(a+b)(a-b)\] so \[(x+8)(x-8)=0\] \[x=8, -8\]

OpenStudy (anonymous):

so kind of like absolute values?

OpenStudy (ash2326):

Yes, so the domain would be all real no.s except (8, -8)

OpenStudy (ash2326):

Do you understand?

OpenStudy (anonymous):

yes thank you for help :)